Twelve different video games showing substance use were observed and the duration of times of game play​ (in seconds) are listed below. The design of the study justifies the assumption that the sample can be treated as a simple random sample. Use the sample data to construct a 90​% confidence interval estimate of sigma​, the standard deviation of the duration times of game play. Assume that this sample was obtained from a population with a normal distribution. 4,975 3,914 4,103 4,107 4,835 4,653 4,887 4,469 3,903 4,397 4,778 4,701
Answer #1

Here s = 386.794 and n = 12
df = 12 - 1 = 11

α = 1 - 0.9 = 0.1
The critical values for α = 0.1 and df = 11 are ꭓ^2(1-α/2,n-1) = 4.5748 and ꭓ^2(α/2,n-1) = 19.6751

(n 1)s2 (n 1)s 2 Xa/2 2 2 VI

CI = (11*386.794^2/19.6751 , 11*386.794^2/4.5748)
CI = (83644.0772 , 359732.7933)